OpenStudy (anonymous):

Select the x-coordinate of the vertex of the parabola defined by the function f(x) = -7x^2 + 9x + 5.

OpenStudy (anonymous):

well the vertex coord are: \[V(\frac{ -b }{ 2a },\frac{ -\Delta }{ 4a })\] so the "x" coord is:\[\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

your equation is: \[-7x^2+9x+5\] where a=-7; b=9; c=5

OpenStudy (anonymous):

Do you understand

OpenStudy (anonymous):

kinda... this one has me a bit perplexed

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

because i typed the equation in incorrectly

OpenStudy (anonymous):

Select the x-coordinate of the vertex of the parabola defined by the function f(x) = -7x^2 + 9x + 5.

OpenStudy (anonymous):

It was supposed to be to the second power and I believe this is were I became screwed up

OpenStudy (anonymous):

yeah don't worry it happens

OpenStudy (anonymous):

thanks